# Broken Symmetry of Stock Returns - a Modified Jones-Faddy Skew t-Distribution

### v1

@@ 1 Introduction
implying that stochastic variance - and hence volatility - tends to revert to its mean value, $\left<v_{t}\right>=\theta$. One of the important implications of the latter is that for returns accumulated over $\tau$ days, $\mathrm{d}t=\tau$, average variance of returns grows linearly with $\tau$, $\left<dx^{2}\right>=\theta\tau$. Since we are not concerned here with quantities such as leverage [perello2003stochastic, dashti2021distributions] and study distribution of returns, in what follows we will neglect correlations between $\mathrm{d}W^{(1)}$ and $\mathrm{d}W^{(2)}$ [dragulescu2002probability, liu2019distributions] and will largely concentrate on daily returns $\tau=1$.
Numerous models exist for $g(v_{t})$, such as Cox-Ingersoll-Ross [cox1985theory, heston1993closed, dragulescu2002probability, liu2019distributions], multiplicative [praetz1972distribution, nelson1990arch, fuentes2009universal, liu2019distributions], and the combination of the two [dashti2021combined]. Here we will concentrate on multiplicative model since it is the simplest model that predicts power-law tails of the distribution of returns and is the easiest to handle analytically. While power-law tails in returns are not universally agreed upon, there is a strong case for them at least for daily returns, while for accumulated returns power law may persist for a large portion of the tail (see e.g. [farahani2025asymmetry] and below).
@@
which yields an Inverse Gamma [wolfram2025inverse] steady-state distribution (probability density function) for variance
@@
which, in turn, translates into the following distribution for volatility [liu2019distributions]
@@
From eqs. (1) and (5) the distribution of stock returns can be found as a product distribution [ma2014model] of inverse Gamma and normal distribution and the result is a Student t-distribution [praetz1972distribution, fuentes2009universal, liu2019distributions]
@@
Clearly this distribution is even  [^2] In this particular case it is a consequence of normal distribution being even in the product distribution in (1). It is also obvious that the product distribution will be even regardless of a specific form of $g(v_{t}))$ and thus establishes symmetry between gains and losses. This, of course, also applies to the power-law tails, whose exponent is $\left(\frac{2\alpha}{\theta}+3\right)$. However this symmetry is clearly broken for actual data. To wit, the distribution of S&P500 returns has [farahani2025asymmetry]:
@@
While still having SDE underpinning “half Student-t” is obviously not an organic distribution. Additionally, it predicts a negative mean contrary to the empirical evidence above. Consequently, we adopted yet another approach based on a skew extension of Student t-distribution by Jones and Faddy. [jones2001skew, jones2003skew]. Unfortunately, modified Jones-Faddy (mJF) distributions that we use are not cleanly derived from SDE formalism but on the other hand they are close in spirit and yield good fits to the full distribution of returns.
@@ 2 Analytical Framework
to characterize the skew of the distribution, where $\widetilde{m}$ is the median, which is evaluated numerically.  [^3] The standard Fisher-Pearson coefficient can not be analytically obtained in this case since the third moment diverges due to slow power-law decays of the tails [farahani2025asymmetry]. We will consider CDF appropriate for gains and losses separately per, respectively,
$F_{g}(x)=\int_{-\infty}^{x}f(y)\mathrm{d}y$ (9)
and
$F_{l}(x)=\int_{x}^{\infty}f(y)\mathrm{d}y$ (10)
@@ 2.1 Student t-Distribution
where $I(x;a,b)$ is a regularized incomplete beta function [nist2025digital] and, obviously, $m_{1}=\overline{m}=\widetilde{m}=\zeta_{1,2}=0$ and
@@ 2.2 Half Student-t Distribution
where $w_{g}+w_{l}=1$ and $w_{g}/w_{l}$ is the ratio of points under gains and losses [farahani2025asymmetry]. Generally speaking a mixture distribution is not a preferable venue from a physicist’s point of view since it does not follow form a first–principles model. However $f_{g\_hSt}$ and $f_{l\_hSt}$ marginally do.
@@
where $\Gamma(x)$ is a Gamma function [nist2025digital]. Clearly, $\overline{m}=0$ in this model so the sign of the skew $\zeta_{1}$ will be that of $m_{1}$. Another observation is that the number of parameters in half Student-t, aside from $\tau$, is double that of Student t-distribution. One possible simplification of this model is to assume that the mean stochastic volatility governing gains and losses is the same, $\theta_{g}=\theta_{l}=\theta$ so that the difference between gains and losses, including power-law exponents, reduces solely to difference between $\alpha_{g}$ and $\alpha_{l}$.
@@ 2.3 Modified Jones-Faddy Distribution mJF1
mJF1 is a direct descendent of the distribution (7) with one minor and one significant variation. First, as in standard Student distribution [wolfram2025student], a location parameter $\mu$ can be introduced and is introduced here. Obviously it does not affect (1) since the variable can always be shifted by a constant. The second variation introduces a skew (skew t distribution [jones2001skew, jones2003skew]), via $\alpha_{g}$ and $\alpha_{l}$ here. In particular, power-law tails scale as $x^{-\left(2\alpha_{g}/\theta+3\right)}$ at $\infty$ and $x^{-\left(2\alpha_{g}/\theta+3\right)}$ at -$\infty$. This brakes a construct based on (1) and (2) which treats volatility of gains and losses uniformly: substituition $\alpha_{g}=\alpha_{l}=\alpha$ in (20) leads back to (7) (with non-zero location parameter $\mu$). At this point we are unaware of an SDE-based formulation which would result in a distribution (20).
@@
$m_{1}=\mu+\sqrt{\frac{(\alpha_{g}+\alpha_{l})\tau}{(\alpha_{g}/\theta+1)+(\alpha_{l}/\theta+1)}}\frac{\left(\frac{\alpha_{l}}{\theta}-\frac{\alpha_{g}}{\theta}\right)\sqrt{\frac{\alpha_{g}+\alpha_{l}}{\theta}+2}}{2\pi\,\mathrm{B}\!\left(\frac{\alpha_{g}}{\theta}+\frac{1}{2},\frac{1}{2}\right)\mathrm{B}\!\left(\frac{\alpha_{l}}{\theta}+\frac{1}{2},\frac{1}{2}\right)}\\$ (24)
@@
$\overline{m}=\mu-\sqrt{\frac{r^{2}}{1-r^{2}}(\alpha_{g}+\alpha_{l})\tau},\qquad r=\frac{\alpha_{l}-\alpha_{g}}{\alpha_{g}+\alpha_{l}+3\theta}.$ (26)
@@ 2.4 Modified Jones-Faddy Distribution mJF2
$f(x)=C\left(1-\frac{x-\mu}{\sqrt{(z-\mu)^{2}+(\alpha_{g}+\alpha_{l})\tau}}\right)^{\frac{\alpha_{g}}{\theta_{g}}+\frac{3}{2}}\left(1+\frac{x-\mu}{\sqrt{(x-\mu)^{2}+(\alpha_{g}+\alpha_{l})\tau}}\right)^{\frac{\alpha_{l}}{\theta_{l}}+\frac{3}{2}}$ (27)
@@
$m_{1}=\mu+\sqrt{\frac{(\alpha_{g}+\alpha_{l})\tau}{(\alpha_{g}/\theta_{g}+1)+(\alpha_{l}/\theta_{l}+1)}}\frac{\left(\frac{\alpha_{l}}{\theta_{l}}-\frac{\alpha_{g}}{\theta_{g}}\right)\sqrt{\frac{\alpha_{g}}{\theta_{g}}+\frac{\alpha_{l}}{\theta_{l}}+2}}{2\pi\,\mathrm{B}\!\left(\frac{\alpha_{g}}{\theta_{g}}+\frac{1}{2},\frac{1}{2}\right)\mathrm{B}\!\left(\frac{\alpha_{l}}{\theta_{l}}+\frac{1}{2},\frac{1}{2}\right)}\\$ (31)
@@
$\overline{m}=\mu-\sqrt{(\alpha_{g}+\alpha_{l})\tau}\,\frac{\frac{\alpha_{l}}{\theta_{l}}-\frac{\alpha_{g}}{\theta_{g}}}{2\sqrt{\left(\frac{\alpha_{g}}{\theta_{g}}+\frac{3}{2}\right)\left(\frac{\alpha_{l}}{\theta_{l}}+\frac{3}{2}\right)}}.$ (33)
@@ 3 Numerical Results
mJF1 $1.412\times 10^{-4}$ - - - $7.873\times 10^{-5}$ $6.371\times 10^{-5}$ $1.168\times 10^{-3}$
mJF2 - $1.187\times 10^{-4}$ $1.782\times 10^{-4}$ - $6.386\times 10^{-5}$ $7.634\times 10^{-5}$ $1.142\times 10^{-3}$
@@
mJF1 $4.57\times 10^{-5}$ $1.45\times 10^{-4}$ $5.22\times 10^{-4}$ $3.96\times 10^{-2}$ $3.211\times 10^{-4}$ $2.29\times 10^{-2}$ -3.12 -2.90
mJF2 $3.7\times 10^{-5}$ $1.3\times 10^{-4}$ $5.49\times 10^{-4}$ $4.49\times 10^{-2}$ $3.395\times 10^{-4}$ $2.65\times 10^{-2}$ -3.07 -2.76
@@
Figs. 12-19 demonstrate the results of statistical tests meant to probe goodness of fit. Figs. 12 and 14 show confidence intervals of linear fits and Figs. 16 and 18 show confidence intervals of mJF1 fits. Confidence intervals are obtained using inversion of the binomial distribution [janczura2012black] and we specifically focused on mJF1 as the most transparent and minimal generalization of Student t-distribution. Figs. 13 and 15 show p-values of order-statistics-based U-test [pisarenko2012robust]  [^4] We are using eq. (14) in [liu2023dragon] for evaluation of p-values. for linear fits and Figs. 17 and 19 show p-values for mJF1 fits. It should be noted that both approaches were developed for detections of outliers, such as Dragon Kings and negative Dragon Kings, in the tails of the distributions. For instance values $p<0.05$ and $p>0.95$ would signal a 95% probability of having, respectively, a Dragon King and a negative Dragon King. In simpler terms, if a data point falls outside the confidence interval and/or if its p-value or (1-p)-value is very small, then it most likely does not belong to the fitting distribution (linear fit being the tail of Pareto distribution).
@@ 4 Conclusions and Discussion
There are a number of possible future directions of this work. The most obvious is to consider other market indices. From our previous experience, we expect that DOW will not exhibit significant difference with S&P500, both in overall behavior and values of parameters. However other long-lasting US indices, such as Russel and NASDAQ, are worth looking into, as well as European and Asian ones. Also overall market returns, reflected by key indices, versus individual companies is a rather challenging question [albuquerque2012skewnwss]. Yet another important avenue is the study of accumulated returns versus dailies: most importantly realized volatility, which shows linear behavior as the function of the number of days of accumulation, and rather abrupt drop off in the tails for longer accumulations. Of course, the most challenging task is finding first-principles explanation of symmetry breaking described in this work.
@@ 6 Acknoledgments
We used MathWorks Matlab for numerical work and Wolfram Mathematica for analytical calculations.

### v2

@@ 1 Introduction
implying that stochastic variance - and hence volatility - tends to revert to its mean value, $\left<v_{t}\right>=\theta$. One of the important implications of the latter is that for returns accumulated over $\tau$ days, $\mathrm{d}t=\tau$, average variance of returns grows linearly with $\tau$, $\left<dx^{2}\right>=\theta\tau$. Since we are not concerned here with quantities such as leverage [18, 3] and study distribution of returns, in what follows we will neglect correlations between $\mathrm{d}W^{(1)}$ and $\mathrm{d}W^{(2)}$ [5, 14] and will largely concentrate on daily returns $\tau=1$.
Numerous models exist for $g(v_{t})$, such as Cox-Ingersoll-Ross [2, 8, 5, 14], multiplicative [20, 16, 7, 14], and the combination of the two [4]. Here we will concentrate on multiplicative model since it is the simplest model that predicts power-law tails of the distribution of returns and is the easiest to handle analytically. While power-law tails in returns are not universally agreed upon, there is a strong case for them at least for daily returns, while for accumulated returns power law may persist for a large portion of the tail (see e.g. [6] and below).
@@
which yields an Inverse Gamma [9] steady-state distribution (probability density function) for variance
@@
which, in turn, translates into the following distribution for volatility [14]
@@
From eqs. (1) and (5) the distribution of stock returns can be found as a product distribution [15] of inverse Gamma and normal distribution and the result is a Student t-distribution [20, 7, 14]
@@
Clearly this distribution is even  [^2] In this particular case it is a consequence of normal distribution being even in the product distribution in (1). It is also obvious that the product distribution will be even regardless of a specific form of $g(v_{t}))$ and thus establishes symmetry between gains and losses. This, of course, also applies to the power-law tails, whose exponent is $\left(\frac{2\alpha}{\theta}+3\right)$. However this symmetry is clearly broken for actual data. To wit, the distribution of S&P500 returns has [6]:
@@
While still having SDE underpinning “half Student-t” is obviously not an organic distribution. Additionally, it predicts a negative mean contrary to the empirical evidence above. Consequently, we adopted yet another approach based on a skew extension of Student t-distribution by Jones and Faddy. [12, 11]. Unfortunately, modified Jones-Faddy (mJF) distributions that we use are not cleanly derived from SDE formalism but on the other hand they are close in spirit and yield good fits to the full distribution of returns.
@@ 2 Analytical Framework
to characterize the skew of the distribution, where $\widetilde{m}$ is the median, which is evaluated numerically.  [^3] See Appendix for an alternative derivation of the median. [^4] The third moment and Fisher-Pearson coefficient diverge due to slow power-law decays of the tails per Table 2 and [6]. We will consider CDF appropriate for gains and losses separately per, respectively,
$F_{g}(x)=\int_{-\infty}^{x}f(y)\mathrm{d}y\quad\text{and}\quad F_{l}(x)=\int_{x}^{\infty}f(y)\mathrm{d}y$ (9)
@@ 2.1 Student t-Distribution
where $I(x;a,b)$ is a regularized incomplete beta function [17] and, obviously, $m_{1}=\overline{m}=\widetilde{m}=\zeta_{1,2}=0$ and
@@ 2.2 Half Student-t Distribution
where $w_{g}+w_{l}=1$ and $w_{g}/w_{l}$ is the ratio of points under gains and losses [6]. Generally speaking a mixture distribution is not a preferable venue from a physicist’s point of view since it does not follow form a first–principles model. However $f_{g\_hSt}$ and $f_{l\_hSt}$ marginally do.
@@
where $\Gamma(x)$ is a Gamma function [17]. Clearly, $\overline{m}=0$ in this model so the sign of the skew $\zeta_{1}$ will be that of $m_{1}$. Another observation is that the number of parameters in half Student-t, aside from $\tau$, is double that of Student t-distribution. One possible simplification of this model is to assume that the mean stochastic volatility governing gains and losses is the same, $\theta_{g}=\theta_{l}=\theta$ so that the difference between gains and losses, including power-law exponents, reduces solely to difference between $\alpha_{g}$ and $\alpha_{l}$.
@@ 2.3 Modified Jones-Faddy Distribution mJF1
mJF1 is a direct descendent of the distribution (7) with one minor and one significant variation. First, as in standard Student distribution [21], a location parameter $\mu$ can be introduced and is introduced here. Obviously it does not affect (1) since the variable can always be shifted by a constant. The second variation introduces a skew (skew t distribution [12, 11]), via $\alpha_{g}$ and $\alpha_{l}$ here. In particular, power-law tails scale as $x^{-\left(2\alpha_{g}/\theta+3\right)}$ at $\infty$ and $x^{-\left(2\alpha_{l}/\theta+3\right)}$ at -$\infty$. This brakes a construct based on (1) and (2) which treats volatility of gains and losses uniformly: substituition $\alpha_{g}=\alpha_{l}=\alpha$ in (19) leads back to (7) (with non-zero location parameter $\mu$). At this point we are unaware of an SDE-based formulation which would result in a distribution (19).
@@
$m_{1}=\mu+\sqrt{(\alpha_{g}+\alpha_{l})\tau}\,\mathrm{B}\!\left(\frac{\alpha_{g}}{\theta}+\frac{1}{2},\frac{1}{2}\right)\,\mathrm{B}\!\left(\frac{\alpha_{l}}{\theta}+\frac{1}{2},\frac{1}{2}\right)\,\frac{\dfrac{\alpha_{l}}{\theta}-\dfrac{\alpha_{g}}{\theta}}{2\pi}\\$ (23)
@@
$\overline{m}=\mu+\sqrt{(\alpha_{g}+\alpha_{l})\tau}\frac{\dfrac{\alpha_{l}}{\theta}-\dfrac{\alpha_{g}}{\theta}}{2\sqrt{\left(\dfrac{\alpha_{g}}{\theta}+\dfrac{3}{2}\right)\left(\dfrac{\alpha_{l}}{\theta}+\dfrac{3}{2}\right)}},$ (25)
@@ 2.4 Modified Jones-Faddy Distribution mJF2
$f(x)=C\left(1-\frac{x-\mu}{\sqrt{(x-\mu)^{2}+(\alpha_{g}+\alpha_{l})\tau}}\right)^{\frac{\alpha_{g}}{\theta_{g}}+\frac{3}{2}}\left(1+\frac{x-\mu}{\sqrt{(x-\mu)^{2}+(\alpha_{g}+\alpha_{l})\tau}}\right)^{\frac{\alpha_{l}}{\theta_{l}}+\frac{3}{2}}$ (26)
@@
$m_{1}=\mu+\sqrt{(\alpha_{g}+\alpha_{l})\tau}\,\mathrm{B}\!\left(\frac{\alpha_{g}}{\theta_{g}}+\frac{1}{2},\frac{1}{2}\right)\,\mathrm{B}\!\left(\frac{\alpha_{l}}{\theta_{l}}+\frac{1}{2},\frac{1}{2}\right)\,\frac{\dfrac{\alpha_{l}}{\theta_{l}}-\dfrac{\alpha_{g}}{\theta_{g}}}{2\pi}\\$ (30)
@@
$\overline{m}=\mu+\sqrt{(\alpha_{g}+\alpha_{l})\tau}\,\frac{\dfrac{\alpha_{l}}{\theta_{l}}-\dfrac{\alpha_{g}}{\theta_{g}}}{2\sqrt{\left(\dfrac{\alpha_{g}}{\theta_{g}}+\dfrac{3}{2}\right)\left(\dfrac{\alpha_{l}}{\theta_{l}}+\dfrac{3}{2}\right)}}.$ (32)
@@ 3 Numerical Results
mJF1 $1.422\times 10^{-4}$ - - - $7.924\times 10^{-5}$ $6.416\times 10^{-5}$ $8.465\times 10^{-4}$
mJF2 - $1.197\times 10^{-4}$ $1.778\times 10^{-4}$ - $6.393\times 10^{-5}$ $7.634\times 10^{-5}$ $8.623\times 10^{-4}$
@@
mJF1 $4.06\times 10^{-5}$ $1.44\times 10^{-4}$ $5.22\times 10^{-4}$ $3.96\times 10^{-2}$ $3.211\times 10^{-4}$ $2.29\times 10^{-2}$ -3.12 -2.90
mJF2 $5.29\times 10^{-5}$ $1.45\times 10^{-4}$ $5.49\times 10^{-4}$ $4.49\times 10^{-2}$ $3.395\times 10^{-4}$ $2.65\times 10^{-2}$ -3.07 -2.76
@@
Figs. 12-19 demonstrate the results of statistical tests meant to probe goodness of fit. Figs. 12 and 14 show confidence intervals of linear fits and Figs. 16 and 18 show confidence intervals of mJF1 fits. Confidence intervals are obtained using inversion of the binomial distribution [10] and we specifically focused on mJF1 as the most transparent and minimal generalization of Student t-distribution. Figs. 13 and 15 show p-values of order-statistics-based U-test [19]  [^5] We are using eq. (14) in [13] for evaluation of p-values. for linear fits and Figs. 17 and 19 show p-values for mJF1 fits. It should be noted that both approaches were developed for detections of outliers, such as Dragon Kings and negative Dragon Kings, in the tails of the distributions. For instance values $p<0.05$ and $p>0.95$ would signal a 95% probability of having, respectively, a Dragon King and a negative Dragon King. In simpler terms, if a data point falls outside the confidence interval and/or if its p-value or (1-p)-value is very small, then it most likely does not belong to the fitting distribution (linear fit being the tail of Pareto distribution).
@@ 4 Conclusions and Discussion
There are a number of possible future directions of this work. The most obvious is to consider other market indices. From our previous experience, we expect that DOW will not exhibit significant difference with S&P500, both in overall behavior and values of parameters. However other long-lasting US indices, such as Russel and NASDAQ, are worth looking into, as well as European and Asian ones. Also overall market returns, reflected by key indices, versus individual companies is a rather challenging question [1]. Yet another important avenue is the study of accumulated returns versus dailies: most importantly realized volatility, which shows linear behavior as the function of the number of days of accumulation, and rather abrupt drop off in the tails for longer accumulations. Of course, the most challenging task is finding first-principles explanation of symmetry breaking described in this work.
@@ 6 Acknoledgments
We used MathWorks Matlab for numerical work and Wolfram Mathematica for analytical calculations. Siqi Shao acknowledges support in part by The University of Cincinnati URC Graduate Support Program.
@@ 7 Conflicts of Interest
Appendix A Derivation of Median
Median is derived from the condition $F(\widetilde{m})=\frac{1}{2}$. For mJF2, for instance, using (28), this gives
$m=\mu+\sqrt{(\alpha_{g}+\alpha_{l})\tau}\;\frac{2u-1}{\sqrt{1-(2u-1)^{2}}}$ (33)
where
$u=I^{-1}\left(\frac{1}{2};\frac{\alpha_{l}}{\theta_{l}}+1,\frac{\alpha_{g}}{\theta_{g}}+1\right)$ (34)
and $I^{-1}$ is the inverse incomplete Beta function. For mJF1 the expression simplifies by setting $\theta_{l}=\theta_{g}=\theta$.